A 5.0 kg mass hangs at the end of a cord. Find the tension in the cord if the acceleration

of the mass is (a)1.5m/s


ଶ downwards (b.) 1.5m/s

upwards


1
Expert's answer
2021-05-10T16:04:28-0400

a)


"T=m(g-a)\\T=5(9.8-1.5)=41.5\ N"

b)


"T=m(g+a)\\T=5(9.8+1.5)=56.5\ N"